# NI-Tex: Non-isometric Image-based Garment Texture Generation

### v1

@@ 1 Introduction
Current texture generation models are typically trained on large-scale 3D datasets such as Objaverse [deitke2023objaverse] and TexVerse [zhang2025texverse], which provide massive collections of 3D assets with PBR textures. These models usually use multi-view attention mechanisms to learn implicit image-geometry correspondences to faithfully transfer the visual appearance and material properties from the image to the target geometry. However, such correspondence learning relies heavily on geometric and topological similarity between the input image and the target geometry. When there exists a significant topological or geometric discrepancy, especially in cases of non-isometric deformation or garment topology variation, the generation quality deteriorates drastically, leading to distorted or inconsistent textures like Figure 2.
Figure 2: Texture generation becomes unreliable when the image prompt and the given mesh exhibit topology inconsistency in Hunyuan3D [hunyuan3d2025hunyuan3d] and Meshy 6 Preview.
In this work, we address this critical challenge by proposing NI-Tex, a training framework that promotes non-isometric image-based garment texture generation, which focuses on robust and controllable texture generation under topological and geometric image-garment inconsistencies. NI-Tex builds upon a texture generation backbone pretrained on Objaverse and TexVerse, and we further fine-tune it using datasets with rich garment deformation data such as BEDLAM [black2023bedlam] to improve the generalization for garment deformations.
@@ 2 Related work
3D Texture Generation methods study how to add textures to given 3D geometries (represented by mesh) from user-defined text or image prompts. TEXTure [richardson2023texture] and Text2Tex [chen2023text2tex] generate images from different viewpoints using diffusion models. These images are then back-projected onto the surface of the geometry. To further improve the alignment between the generated images and the given geometry, Easi-Tex [perla2024easi] uses edge maps to guide the diffusion process. However, images from different viewpoints often cannot fully cover the entire geometry. Thus, some faces of the mesh remain untextured. This problem is more evident for objects with complex structures. Paint3D [zeng2024paint3d] additionally introduces an inpainting step on the UV map of the geometry, allowing all faces to be textured. By incorporating global texture information, it further improves the texture consistency across multiple viewpoints. Although the above methods achieve good results in 3D texture generation, they overlook the process of lighting and shadows in the generated images. As a result, the produced textures fail to deliver realistic visual effects under new lighting conditions.
@@ 2.2 PBR Texture Generation
PBR texture generation methods output the object’s base color (albedo) and its perceptual properties (metallic and roughness) separately. The Score-Distillation-Sampling (SDS) based optimization approach and the data-driven feed-forward approach are two common categories for PBR texture generation. Fantasia3D [chen2023fantasia3d], Matlaber [xu2023matlaber], and Paint-it [youwang2024paint] leverage the priors of pretrained 2D diffusion models and use SDS to distill the PBR texture attributes of the object. FlashTex [deng2024flashtex] and DreamMat [zhang2024dreammat] train light ControlNet to achieve light-aware diffusion models by augmenting the training dataset with simulated reflection effect under various lighting conditions. However, SDS-based optimization is time-consuming and sometimes produces less realistic visual effects. Another line of research focuses on feed-forward training enabled by existing open-source 3D datasets [deitke2023objaverse, deitke2023objaverse-xl, zhang2025texverse]. MaterialAnything [huang2025material] and Hunyuan3D 2.1 [hunyuan3d2025hunyuan3d] render PBR-annotated objects from the datasets (mainly from Objaverse) under different lighting conditions and viewpoints. They use albedo, metallic, and roughness maps for supervision in training, which makes the model material-aware. The feed-forward approach effectively reduces inference time, enabling efficient and large-scale PBR texture generation. Hunyuan3D 2.1 applies rigid operations such as scaling, translation, and rotation during rendering to augment training data for better generalization. However, the model still struggles to generate textures between image-geometry pairs with non-isometric differences such as topology changes.
@@ 2.3 Image-based Virtual Try-Off and Image-Prompt 3D Garment Texture Generation
TryOffDiff [velioglu2024tryoffdiff] is the first to introduce Virtual Try-Off (VTOFF), which extracts garment texture from given photos of clothed individuals and generates standardized 2D garment images. OMFA [liu2025one] and Voost [lee2025voost] build a unified framework that combines Virtual Try-On (VTON) and VTOFF techniques to achieve better results. However, they are limited to generating 2D images and cannot bake the extracted garment textures onto a given 3D geometry. In the field of image-based 3D garment generation, methods such as Pix2Surf [mir2020learning], Cloth2Tex [gao2024cloth2tex], and Garment3DGen [sarafianos2025garment3dgen] typically require a 2D garment image with canonical pose and then align a template mesh to the garment image through mesh deformation, which suffers from compound errors from the mesh alignment procedure and has limited applicability.
@@ 3.2 Network Architecture
We adopt the MCAA module from MaterialMVP [he2025materialmvp] to connect the two branches. As defined in Equation (1), we extract the reference texture features from the guidance branch and inject them into the albedo channel of the training branch:
@@ 4.1 Dataset Preparation
Cross-pose Augmentation. We construct 3D Garment Videos based on the BEDLAM[black2023bedlam] dataset. BEDLAM contains nearly 1,700 distinct garment albedo textures, each associated with hundreds of motion sequences, and each sequence consists of hundreds of frames. Since the original dataset does not provide roughness or metallic attributes, we augment the selected motion sequences by assigning PBR material properties to each frame. Specifically, we set roughness and metallic values as:
@@ 4.2 Iterative Baking via View Selection and Uncertainty Quantification
Uncertainty Quantification. We formulate the Uncertainty Quantification task as an Image Quality Assessment (IQA) problem following Active View Selector (AVS) [wang2025active]. Given an single-view texture map $I\in\mathbb{R}^{H\times W\times 3}$, the UQ model adopts an residual net [he2016deep] architecture and learns to predict per-pixel uncertainty scores $S\in\mathbb{R}^{H\times W}$ on the multi-view texture maps via supervised learning.
@@ 5.1 Experimental Setting
Our diffusion backbone is based on Stable Diffusion 2.1, trained on 8 H200 GPUs for about 10 days with a batch size of 2 and an image resolution of 512×512. For evaluation, we use both industrial well-rendered images and wild images from DeepFashion2 [ge2019deepfashion2] masked by SAM2 [ravi2024sam] as input prompts, and select industrial 3D meshes together with Hunyuan3D-generated meshes as target geometries.
@@ 5.2 Qualitative Evaluation
Effectiveness of Cross-pose Texture Generation. The 4D-Dress dataset [wang20244d] provides per-frame 3D captures of real human motion, though its texture maps often contain baked shadows. We use 4D-Dress to maintain consistent garment topology across poses, enabling a focused evaluation of NI-Tex’s cross-pose texture generation (See Figure 7). For each sequence, we randomly sample a mesh and viewpoint to render the input image prompt, and use another pose of the same subject as the target mesh.

### v2

@@ 1 Introduction
Current texture generation models are typically trained on large-scale 3D datasets such as Objaverse [1] and TexVerse [2], which provide massive collections of 3D assets with PBR textures. These models usually use multi-view attention mechanisms to learn implicit image-geometry correspondences to faithfully transfer the visual appearance and material properties from the image to the target geometry. However, such correspondence learning relies heavily on geometric and topological similarity between the input image and the target geometry. When there exists a significant topological or geometric discrepancy, especially in cases of non-isometric deformation or garment topology variation, the generation quality deteriorates drastically, leading to distorted or inconsistent textures like Figure 2.
Figure 2: Texture generation becomes unreliable when the image prompt and the given mesh exhibit topology inconsistency in Hunyuan3D [3] and Meshy 6 Preview.
In this work, we address this critical challenge by proposing NI-Tex, a training framework that promotes non-isometric image-based garment texture generation, which focuses on robust and controllable texture generation under topological and geometric image-garment inconsistencies. NI-Tex builds upon a texture generation backbone pretrained on Objaverse and TexVerse, and we further fine-tune it using datasets with rich garment deformation data such as BEDLAM [4] to improve the generalization for garment deformations.
@@ 2 Related work
3D Texture Generation methods study how to add textures to given 3D geometries (represented by mesh) from user-defined text or image prompts. TEXTure [5] and Text2Tex [6] generate images from different viewpoints using diffusion models. These images are then back-projected onto the surface of the geometry. To further improve the alignment between the generated images and the given geometry, Easi-Tex [7] uses edge maps to guide the diffusion process. However, images from different viewpoints often cannot fully cover the entire geometry. Thus, some faces of the mesh remain untextured. This problem is more evident for objects with complex structures. Paint3D [8] additionally introduces an inpainting step on the UV map of the geometry, allowing all faces to be textured. By incorporating global texture information, it further improves the texture consistency across multiple viewpoints. Although the above methods achieve good results in 3D texture generation, they overlook the process of lighting and shadows in the generated images. As a result, the produced textures fail to deliver realistic visual effects under new lighting conditions.
@@ 2.2 PBR Texture Generation
PBR texture generation methods output the object’s base color (albedo) and its perceptual properties (metallic and roughness) separately. The Score-Distillation-Sampling (SDS) based optimization approach and the data-driven feed-forward approach are two common categories for PBR texture generation. Fantasia3D [9], Matlaber [10], and Paint-it [11] leverage the priors of pretrained 2D diffusion models and use SDS to distill the PBR texture attributes of the object. FlashTex [12] and DreamMat [13] train light ControlNet to achieve light-aware diffusion models by augmenting the training dataset with simulated reflection effect under various lighting conditions. However, SDS-based optimization is time-consuming and sometimes produces less realistic visual effects. Another line of research focuses on feed-forward training enabled by existing open-source 3D datasets [1, 14, 2]. MaterialAnything [15] and Hunyuan3D 2.1 [3] render PBR-annotated objects from the datasets (mainly from Objaverse) under different lighting conditions and viewpoints. They use albedo, metallic, and roughness maps for supervision in training, which makes the model material-aware. The feed-forward approach effectively reduces inference time, enabling efficient and large-scale PBR texture generation. Hunyuan3D 2.1 applies rigid operations such as scaling, translation, and rotation during rendering to augment training data for better generalization. However, the model still struggles to generate textures between image-geometry pairs with non-isometric differences such as topology changes.
@@ 2.3 Image-based Virtual Try-Off and Image-Prompt 3D Garment Texture Generation
TryOffDiff [16] is the first to introduce Virtual Try-Off (VTOFF), which extracts garment texture from given photos of clothed individuals and generates standardized 2D garment images. OMFA [17] and Voost [18] build a unified framework that combines Virtual Try-On (VTON) and VTOFF techniques to achieve better results. However, they are limited to generating 2D images and cannot bake the extracted garment textures onto a given 3D geometry. In the field of image-based 3D garment generation, methods such as Pix2Surf [19], Cloth2Tex [20], and Garment3DGen [21] typically require a 2D garment image with canonical pose and then align a template mesh to the garment image through mesh deformation, which suffers from compound errors from the mesh alignment procedure and has limited applicability.
@@ 3.2 Network Architecture
We adopt the MCAA module from MaterialMVP [22] to connect the two branches. As defined in Equation (1), we extract the reference texture features from the guidance branch and inject them into the albedo channel of the training branch:
@@ 4.1 Dataset Preparation
Cross-pose Augmentation. We construct 3D Garment Videos based on the BEDLAM[4] dataset. BEDLAM contains nearly 1,700 distinct garment albedo textures, each associated with hundreds of motion sequences, and each sequence consists of hundreds of frames. Since the original dataset does not provide roughness or metallic attributes, we augment the selected motion sequences by assigning PBR material properties to each frame. Specifically, we set roughness and metallic values as:
@@ 4.2 Iterative Baking via View Selection and Uncertainty Quantification
Uncertainty Quantification. We formulate the Uncertainty Quantification task as an Image Quality Assessment (IQA) problem following Active View Selector (AVS) [23]. Given an single-view texture map $I\in\mathbb{R}^{H\times W\times 3}$, the UQ model adopts an residual net [24] architecture and learns to predict per-pixel uncertainty scores $S\in\mathbb{R}^{H\times W}$ on the multi-view texture maps via supervised learning.
@@ 5.1 Experimental Setting
Our diffusion backbone is based on Stable Diffusion 2.1, trained on 8 H200 GPUs for about 10 days with a batch size of 2 and an image resolution of 512×512. For evaluation, we use both industrial well-rendered images and wild images from DeepFashion2 [25] masked by SAM2 [26] as input prompts, and select industrial 3D meshes together with Hunyuan3D-generated meshes as target geometries.
@@ 5.2 Qualitative Evaluation
Effectiveness of Cross-pose Texture Generation. The 4D-Dress dataset [27] provides per-frame 3D captures of real human motion, though its texture maps often contain baked shadows. We use 4D-Dress to maintain consistent garment topology across poses, enabling a focused evaluation of NI-Tex’s cross-pose texture generation (See Figure 7). For each sequence, we randomly sample a mesh and viewpoint to render the input image prompt, and use another pose of the same subject as the target mesh.